OpenStudy (anonymous):

t 96 ____ = ____ Proportion help, pwease? :3 6 16

OpenStudy (anonymous):

t/6 and 96/16

OpenStudy (anonymous):

To find t, you need to cross multiply. This works with all proportions. a/b=c/d and also da=cb In regard to your equation, 16t=96(6) Simplify and then divide both sides by 16 to get your answer.

OpenStudy (anonymous):

Thank your, Mister Justin, sir! :D